Car on curve 12 3 4 5 67 A car is traveling around a horizontal circular track with radius r = 240 m at a constant speed v angle 8A 29 above the x axis, and the angle Op 58 below the x axis. 25 m/s as shown. The 1) What is the magnitude of the cars acceleration? m/s Submit 2) What is the x component of the cars acceleration when it is at point A m/ st Submit 3) What is the y component of the cars acceleration when it is at point A sSubmit
m/s Submit 3) What is the y component of the cars acceleration when it is at point A m/s Submit 4) What is the x component of the cars acceleration when it is at point B m/s2 Submit 5) What is the y component of the cars acceleration when it is at point B m/ s2 Submit 6) As the car passes point B, the y component of its acceleration is O increasing constant O decreasing Submit
Answer #1

1)

acceleration = v2/r = (25)2/(240) = 2.60 m/s2

2)

x component of acceleration is

ax = (2.60)*(cos 29) = 2.27 m/s2 (In the negative x direction)

3)

y component of acceleration is

ay = 2.60*(sin 29) = 1.26 m/s2

4)

x component = (2.60)(cos 58) = 1.37 m/s2 (In the negative x direction )

5)

y component = (2.60)(sin 58) = 2.20 m/s2

6)

As the car passes the point B, the angle \theta b decreases, and sin\thetab decreases.

So, the y-component of acceleration ( a sin\thetab) decreases.